OpenStudy (anonymous):

Solve the pair of simultaneous equations. 3x-2y=10 2x+5y=13

OpenStudy (anonymous):

hmm lets see

OpenStudy (anonymous):

first solve the top equation for y. Can you do this?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

can you try?

OpenStudy (anonymous):

i dnt know what to do @Outkast3r09

OpenStudy (anonymous):

http://www.youtube.com/watch?v=cwHR_B9zK7k

OpenStudy (anonymous):

can u make y the subject of the formula? @Tashagaye

OpenStudy (anonymous):

of which formula @hubertH

OpenStudy (anonymous):

@Outkast3r09 i dnt understan why cant u help me i dnt understan

OpenStudy (anonymous):

the question

OpenStudy (anonymous):

Well first you need to go back and review the properties or equalities before you try to handle this problem

OpenStudy (anonymous):

These properties simply state that basically, if you use a sign on one side of an equality, you must do the same sign on the right side to keep them equal to eachother (this includes subtraction, addition, multiplication, division, exponents, radicals)

OpenStudy (anonymous):

Addition \[x=y\] \[2=2\] \[x+2\neq y\] \[2+2\neq 2\] \[x+2=y+2\] \[4=4\]

OpenStudy (anonymous):

3x - 2y =10 so minus 10 both sides 3x-2y-10=0 so add 2y both sides 3x-10=2y divide both sides by 2 3x-10/2=y

OpenStudy (anonymous):

hubert, especially don't give the answer in this case... properties of equalities are needed for almost everything

OpenStudy (anonymous):

i still dnt get ur working out @Outkast3r09

OpenStudy (anonymous):

alright so if you have something such sa \[x=y\] in which a set such as (2,2) works, if you manipulate one side of the equation, the equality no longer stands. See how when i added two to one side and not the other, the equality broke and it was no longer equal. However when i added two to both sides it kept it equal

OpenStudy (anonymous):

so using , addition subtraction division and multiplication, you want to get \[y=something\] for the first equation